# Rethinking Graph Structure Learning in the Era of LLMs

### v1

@@ D.3 Leaf-oriented Two-step Sampling
Input: Original Graph $\mathcal{G}$, Optimized encoding tree $\mathcal{T}^{\star}$, soft label $\mathbf{Y}^{cls}$ for each leaf node, hyperparameters $\theta$, $r$.
Output: Optimized graph $\mathcal{G}^{*}$.
for $\mathcal{C}_{l}\in\mathcal{T}^{\star}$ do
       for $i=1$ to $m\times r$ do
             // Step 1: Sample a leaf node $\alpha$ from $\mathcal{C}_{l}$
Let $set_{1}=\{\alpha|\alpha\in\mathcal{C}_{l}\}$;
for $\alpha\in set_{1}$ do
                   Compute $\mathcal{H}^{\mathcal{T}^{\star}}(\mathcal{G},\alpha)$ according to Eq. (10);
Compute $\mathcal{P}_{topo}(\alpha)$ according to Eq. (7);
            Sample a leaf node $\alpha$ based on $\mathcal{P}_{topo}(\alpha)$ from $set_{1}$;
// Step 2: Sample an edge for $\alpha$
Let $set_{2}=\{\beta|\beta\in\mathcal{C}_{l}$ and $\beta\neq\alpha\}$;
if $|set_{2}|<\theta$ then
                   // Expand the candidate set
Let $\delta$ be the grandparent of $\alpha$;
for $\gamma\in\delta$ and $\gamma\neq\alpha^{+}$ do
                         Compute $y_{\gamma}^{cls}=\frac{1}{m}\sum_{\alpha\in\gamma}y_{\alpha}^{cls}$;
Compute $\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha})$ according to Eq. (3);
                  for $|set_{2}|<\theta$ do
                         Choose low-level community with highest $\mathrm{sim}$ and add its children to $set_{2}$;
            for $\beta\in set_{2}$ do
                   Compute $\mathrm{sim}(y^{cls}_{\beta},y^{cls}_{\alpha})$ according to Eq. (3);
Compute $\mathcal{P}_{sema}^{\alpha}(\beta)$ according to Eq. (8);
            if $|set_{2}|>\theta$ then
                   Keep the top $\theta$ nodes in $set_{2}$ based on the soft label similarity with $\alpha$.
            For edge addition, sample a leaf node $\beta$ from $set_{2}$ based on $\mathcal{P}_{sema}^{\alpha}(\beta)$ ranked in descending order;
For edge removal, sample a leaf node $\beta$ from $set_{2}$ based on $\mathcal{P}_{sema}^{\alpha}(\beta)$ ranked in ascending order;
Add edge to or remove edge from graph $\mathcal{G}$;
Return $\mathcal{G}^{\star}\leftarrow\mathcal{G}$;
Algorithm 3 Leaf-oriented Two-step Sampling
Appendix E Datasets
Our framework LLaTA is evaluated on 10 TAG datasets, the details of these TAG datasets are shown in Table 5. The description of the datasets for each domain is as follows:
@@ Appendix E Datasets
  • Citation Networks  (Chen et al., 2024). Cora, Citeseer, and Pubmed are benchmark datasets of citation networks. Nodes represent paper, and edges represent citation relationships. The features are obtained through pre-trained language model, and labels denote their academic fields.
@@
  • Knowledge Networks  (Chen et al., 2024; Mernyei & Cangea, 2020). WikiCS is a benchmark dataset of knowledge networks. Nodes represent articles in the field of computer science, and edges represent hyperlinks between these articles. The features are obtained through pre-trained language model, and labels denote different branches of computer science.
@@
  • Social Networks  (Huang et al., 2024). Instagram and Reddit are benchmark datasets of social netowrks. Nodes represent users, and edges represent social relationhships. The features are obtained through pre-trained language model, and labels correspond to different types of users.
@@
  • E-commerce Networks  (Yan et al., 2023). Ratings, Child, History and Photo are benchmark datasets of e-commerce networks. Nodes represent products, and edges represent relationships such as co-purchase or co-view. The features are obtained through pre-trained language model, and labels correspond to product categories or user ratings.
Appendix F Baselines
@@
In this section, we provide a brief description for each baseline used in the experiments. For LLM-based methods, we primarily select GLM-4-9B as the LLM backbone. For LLaTA, we utilize the GLM-4 to perform inference on the node classification probabilities. For LLM4RGNN, we use the GLM-4 to construct the fine-tuning dataset and fine-tune the Llama-3-8B for edge prediction; for GraphEdit, we use the GLM-4 for both fine-tuning and edge prediction.
@@
  • GCN (Kipf & Welling, 2016). GCN is among the most widely adopted GNN architectures, as it introduces a first-order approximation of a localized spectral filter for graph-structured data.
@@
  • GAT (Veličković et al., 2017). GAT incorporates a self-attention mechanism to compute importance scores for different neighboring nodes, enabling more effective aggregation of neighborhood information.
@@
  • GraphSAGE (Hamilton et al., 2017). GraphSAGE [14] is an inductive framework that generates node embeddings by leveraging node features and using a sampling-based approach to aggregate information from the local neighborhood, enabling scalable representation learning.
@@
  • IDGL (Chen et al., 2020). IDGL is a framework that learns a refined graph structure by leveraging cosine similarity of features and top-k threshold refinement, combining it with the original graph for joint optimization.
@@
  • SLAPS (Fatemi et al., 2021). SLAPS is a graph learning framework that uses MLP-based embeddings and kNN to build the graph. The adjacency matrix is symmetrized and normalized, and a self-supervised denoising autoencoder is used to update both the graph and model parameters.
@@
  • GAUGO (Zhao et al., 2021). GAUGO uses a graph auto-encoder to learn a new graph structure. It predicts edges based on node features, refines the edge probabilities with Gumbel sampling, and fuses the generated graph with the original one before training. Both the generated graph and model parameters are updated during optimization.
@@
  • HESGSL (Wu et al., 2023). HESGSL uses hierarchical embeddings to capture both local and global graph structures. It employs self-supervised learning with clustering and attention mechanisms to enhance feature representation and performance on downstream tasks.
@@
  • SEGSL (Zou et al., 2023). SEGSL constructs a graph by fusing a kNN graph with the original graph, leveraging structural entropy and encoding tree to refine edges and hierarchically extract community structures. The graph and model parameters are optimized jointly during training.
@@
  • ProGNN (Jin et al., 2020). ProGNN refines graph structures by enforcing low-rank, sparsity, and similarity constraints, jointly optimizing the learned graph and model parameters without predefined GSL bases or graph fusion.
@@
  • SUBLIME (Liu et al., 2022b). SUBLIME constructs GSL graphs using an anchor view (original graph) and a learner view. The learn view is initialized with kNN and optimized with parameterized or non-parameterized methods, followed by post-processing steps like top-k filtering and symmetrization. Contrastive learning between the two views refines the graph for downstream tasks.
@@
  • STABLE (Li et al., 2022). STABLE employs Graph Contrastive Learning (GCL) to generate robust graphs by perturbing node similarities and edges. Positive samples use slight perturbations, while negative samples use shuffled features. The graph refinement step applies a top-k filtering strategy on the node similarity matrix to retain helpful edges while removing adversarial ones.
@@
  • CoGSL (Liu et al., 2022a). CoGSL constructs GSL graphs by combining the original graph with additional views, such as adjacency or kNN graphs. Node embeddings are generated via GCNs, and connection probabilities are estimated and refined using InfoNCE loss to maximize mutual information. The final graph is obtained through early fusion and updated with model parameters.
@@
  • BORF (Nguyen et al., 2023). BORF is a curvature-based graph rewiring approach aimed at addressing over-smoothing and over-squashing in GNNs. It leverages Ollivier-Ricci curvature to identify edges with extreme curvature values—positive for over-smoothing and negative for over-squashing. This process preserves the graph’s topology while improving message-passing efficiency and enhancing GNN performance on downstream tasks.
@@
  • DHGR (Bi et al., 2024). DHGR enhances GNN performance on heterophily graphs by preprocessing the graph structure through rewiring. It systematically adds homophilic edges and removes heterophilic ones based on label and feature similarities, improving node classification accuracy, particularly for heterophily settings.
@@
  • GraphEdit (Guo et al., 2024). GraphEdit refines graph structures by integrating LLMs with a lightweight edge predictor to enhance graph representation learning. By leveraging instruction-tuning and training edge predictor, GraphEdit effectively denoises noisy connections and identifies implicit relationships among nodes.
@@
  • LLM4RGNN (Zhang et al., 2024). LLM4RGNN is a framework that enhances the adversarial robustness of GNNs using LLMs. It detects malicious edges and restores critical ones through a purification strategy, leveraging fine-tuned LLMs for edge prediction and correction. This approach effectively mitigates topology attacks, improving GNN performance across attack scenarios.
Appendix G Case Study
@@
Figure 7: A case study of the tree-prompted LLaTA inference.
@@
Appendix H LLM Backbone Analysis
@@
Appendix I Hyperparameter Analysis
@@
Figure 8: The performance of LLaTA with different $\epsilon$.
@@
To further answer Q3, we analyze the impact of the similarity threshold $\epsilon$ in Sec.4.3. Specifically, we evaluate the node classification performance of LLaTA under varying $\epsilon$ on the Cora and Citeseer, with the results presented in Fig. 8.
@@
The experimental findings indicate that the hyperparameter $\epsilon$ is critical in determining the quality of the in-context information provided to LLMs and the effectiveness of their inference. A small $\epsilon$ value may lead to the aggregation of nodes from different classes, introducing noise into the in-context and thereby reducing inference accuracy. Conversely, an excessively large $\epsilon$ may result in over-filtering, disregarding important connections between similar nodes, ultimately leading to a loss of meaningful relationships and failing to provide high-quality contextual information for LLMs. Our results suggest that an appropriate range for $\epsilon$ is between 0.4 and 0.6, which balances the trade-off between minimizing noise and preserving essential connections. Specifically, the optimal $\epsilon$ value is 0.45 for Cora and 0.55 for Citeseer. In conclusion, selecting an appropriate $\epsilon$ is crucial for constructing high-quality in-context information for LLMs and enhancing the overall performance of LLaTA.

### v2

@@ F.4 Structural Entropy and Node Topological Uncertainty
Appendix G Datasets
We evaluate LLaTA and baselines on 10 widely adopted TAG datasets across multiple domains, including 3 citation networks [3], 1 knowledge network [3], 2 social networks [9], and 4 e-commerce networks [29]. Among these, the Ratings and Child exhibit heterophily, while the remaining datasets follow homophily. The details of these TAG datasets are shown in Table 6. The description of the datasets for each domain is as follows:
@@ Appendix G Datasets
  • Citation Networks  [3]. Cora, Citeseer, and Pubmed are benchmark datasets of citation networks. Nodes represent paper, and edges represent citation relationships. The features are obtained through pre-trained language model, and labels denote their academic fields.
@@
  • Knowledge Networks  [3, 20]. WikiCS is a benchmark dataset of knowledge networks. Nodes represent articles in the field of computer science, and edges represent hyperlinks between these articles. The features are obtained through pre-trained language model, and labels denote different branches of computer science.
@@
  • Social Networks  [9]. Instagram and Reddit are benchmark datasets of social netowrks. Nodes represent users, and edges represent social relationhships. The features are obtained through pre-trained language model, and labels correspond to different types of users.
@@
  • E-commerce Networks  [29]. Ratings, Child, History and Photo are benchmark datasets of e-commerce networks. Nodes represent products, and edges represent relationships such as co-purchase or co-view. The features are obtained through pre-trained language model, and labels correspond to product categories or user ratings.
Appendix H Baselines
@@
In this section, we provide a brief description for each baseline used in the experiments. It includes 5 coupled GSL methods [2, 6, 32, 28, 37], 6 decoupled GSL methods [11, 18, 14, 16, 21, 1], and 2 LLM-based GSL methods[7, 31]. Based on this, we apply prevalent GCN [12], GAT [26], and GraphSAGE [8] as the downstream backbone. For LLM-based methods, we primarily select GLM-4-9B as the LLM backbone. For LLaTA, we utilize the GLM-4 to perform inference on the node classification probabilities. For LLM4RGNN, we use the GLM-4 to construct the fine-tuning dataset and fine-tune the Llama-3-8B for edge prediction; for GraphEdit, we use the GLM-4 for both fine-tuning and edge prediction.
@@
  • GCN [12]. GCN is among the most widely adopted GNN architectures, as it introduces a first-order approximation of a localized spectral filter for graph-structured data.
@@
  • GAT [26]. GAT incorporates a self-attention mechanism to compute importance scores for different neighboring nodes, enabling more effective aggregation of neighborhood information.
@@
  • GraphSAGE [8]. GraphSAGE [14] is an inductive framework that generates node embeddings by leveraging node features and using a sampling-based approach to aggregate information from the local neighborhood, enabling scalable representation learning.
@@
  • IDGL [2]. IDGL is a framework that learns a refined graph structure by leveraging cosine similarity of features and top-k threshold refinement, combining it with the original graph for joint optimization.
@@
  • SLAPS [6]. SLAPS is a graph learning framework that uses MLP-based embeddings and kNN to build the graph. The adjacency matrix is symmetrized and normalized, and a self-supervised denoising autoencoder is used to update both the graph and model parameters.
@@
  • GAUGO [32]. GAUGO uses a graph auto-encoder to learn a new graph structure. It predicts edges based on node features, refines the edge probabilities with Gumbel sampling, and fuses the generated graph with the original one before training. Both the generated graph and model parameters are updated during optimization.
@@
  • HESGSL [28]. HESGSL uses hierarchical embeddings to capture both local and global graph structures. It employs self-supervised learning with clustering and attention mechanisms to enhance feature representation and performance on downstream tasks.
@@
  • SEGSL [37]. SEGSL constructs a graph by fusing a kNN graph with the original graph, leveraging structural entropy and encoding tree to refine edges and hierarchically extract community structures. The graph and model parameters are optimized jointly during training.
@@
  • ProGNN [11]. ProGNN refines graph structures by enforcing low-rank, sparsity, and similarity constraints, jointly optimizing the learned graph and model parameters without predefined GSL bases or graph fusion.
@@
  • SUBLIME [18]. SUBLIME constructs GSL graphs using an anchor view (original graph) and a learner view. The learn view is initialized with kNN and optimized with parameterized or non-parameterized methods, followed by post-processing steps like top-k filtering and symmetrization. Contrastive learning between the two views refines the graph for downstream tasks.
@@
  • STABLE [14]. STABLE employs Graph Contrastive Learning (GCL) to generate robust graphs by perturbing node similarities and edges. Positive samples use slight perturbations, while negative samples use shuffled features. The graph refinement step applies a top-k filtering strategy on the node similarity matrix to retain helpful edges while removing adversarial ones.
@@
  • CoGSL [16]. CoGSL constructs GSL graphs by combining the original graph with additional views, such as adjacency or kNN graphs. Node embeddings are generated via GCNs, and connection probabilities are estimated and refined using InfoNCE loss to maximize mutual information. The final graph is obtained through early fusion and updated with model parameters.
@@
  • BORF [21]. BORF is a curvature-based graph rewiring approach aimed at addressing over-smoothing and over-squashing in GNNs. It leverages Ollivier-Ricci curvature to identify edges with extreme curvature values—positive for over-smoothing and negative for over-squashing. This process preserves the graph’s topology while improving message-passing efficiency and enhancing GNN performance on downstream tasks.
@@
  • DHGR [1]. DHGR enhances GNN performance on heterophily graphs by preprocessing the graph structure through rewiring. It systematically adds homophilic edges and removes heterophilic ones based on label and feature similarities, improving node classification accuracy, particularly for heterophily settings.
@@
  • GraphEdit [7]. GraphEdit refines graph structures by integrating LLMs with a lightweight edge predictor to enhance graph representation learning. By leveraging instruction-tuning and training edge predictor, GraphEdit effectively denoises noisy connections and identifies implicit relationships among nodes.
@@
  • LLM4RGNN [31]. LLM4RGNN is a framework that enhances the adversarial robustness of GNNs using LLMs. It detects malicious edges and restores critical ones through a purification strategy, leveraging fine-tuned LLMs for edge prediction and correction. This approach effectively mitigates topology attacks, improving GNN performance across attack scenarios.
Appendix I Hyperparameter Settings
[…]
For LLaTA, the hyperparameters are searched as follows: The height of encoding tree $K$ in Sec. 4.2 is chosen from {2, 3, 4, 5}, the similarity threshold $\epsilon$ in Sec. 4.3 is selected from [0.4, 0.6], while the size of the candidate set $\theta$ in Sec. 4.4 is searched in {3, 5, 10, 15, 20}, and the two-step sampling frequency $r$ is tuned among {1, 3, 5, 10, 25}.
For other baselines, we primarily adopt the optimal hyperparameters reported in their original papers. If no such configurations are available, we utilize Optuna to automatically determine the best hyperparameter settings for these baselines.
Appendix J Case Study
@@ Appendix I Hyperparameter Settings
Figure 7: A case study of the tree-prompted LLaTA inference.
@@
Appendix K LLM Backbone Analysis
@@
Appendix L Hyperparameter Analysis
@@
Figure 8: The performance of LLaTA with different $\epsilon$ on Cora and Citeseer datasets.
Figure 9: The performance of LLaTA with different $\theta$ and $r$ on Child and History datasets.
To further answer Q3, we analyze the impact of the similarity threshold $\epsilon$ in Sec. 4.3. Additionally, we conducted further analysis on the impact of parameters $\theta$ and $r$ in Sec. 4.4 across more datasets.
L.1 Analysis of Hyperparameter $\epsilon$
@@
Specifically, we evaluate the node classification performance of LLaTA under varying $\epsilon$ on the Cora and Citeseer, with the results presented in Fig. 9.
@@
The experimental findings indicate that the hyperparameter $\epsilon$ is critical in determining the quality of the in-context information provided to LLMs and the effectiveness of their inference. A small $\epsilon$ value may lead to the aggregation of nodes from different classes, introducing noise into the in-context and thereby reducing inference accuracy. Conversely, an excessively large $\epsilon$ may result in over-filtering, disregarding important connections between similar nodes, ultimately leading to a loss of meaningful relationships and failing to provide high-quality contextual information for LLMs. Our results suggest that an appropriate range for $\epsilon$ is between 0.4 and 0.6, which balances the trade-off between minimizing noise and preserving essential connections. Specifically, the optimal $\epsilon$ value is 0.45 for Cora and 0.55 for Citeseer. In conclusion, the selection of an appropriate $\epsilon$ value plays a critical role in constructing high-quality in-context information for LLMs, ultimately enhancing the overall performance and robustness of LLaTA across diverse tasks and datasets.
L.2 Analysis of Hyperparameters $\theta$ and $r$
To thoroughly analyze the impact of $\theta$ and $r$ on LLaTA’s performance, we conducted additional experiments on the Child and History datasets, with the results presented in Fig. 9.
Our evaluation on the Child dataset shows pronounced sensitivity to the two-step sampling frequency $r$. Accuracy peaks at $r=10$ (62.28%) before declining to 61.69% at $r=25$, confirming that excessive edge additions ($r\geq 25$) disrupt heterophilic structures. The semantic sampling parameter $\theta$ maintains stable performance (61.87%-61.98%) within $[5,15]$, while extreme values ($\theta=3$ or $20$) cause 1-2% drops due to sampling limitations. The History dataset demonstrates LLaTA’s robustness to $r$ variations (84.32%-85.28% accuracy), with only 0.02% difference between $r=5$ (85.34%) and $r=25$ (85.32%). For $\theta$, peak accuracy at $\theta=10$ (85.28%) varies by less than 1% across $[5,15]$, This systematic comparison reveals that $\theta$ optimization should be performed after determining $r$, particularly for heterophilic graphs where $r$ significantly impacts structural integrity while $\theta$ mainly affects sampling quality.
These findings collectively validate the dataset-dependent nature of $r$’s impact and the universal stability of $\theta$ observed in our primary experiments. The Child dataset’s heterophily demands careful $r$ calibration to avoid overhomophilization, whereas History’s homophily permits broader parameter flexibility. Both datasets corroborate that $\theta\in[5,15]$ serves as a reliable default range, though practitioners should prioritize $r$ tuning when deploying LLaTA on heterophilic graphs. The consistent performance patterns across three distinct graph types (Cora, Child, History) strengthen the generalizability of our parameter guidelines.
Appendix M Efficiency Analysis of LLaTA
Table 8: Complete runtime(h) results of LLaTA across all 10 datasets.
Method Cora Citeseer Pubmed WikiCS Instagram Reddit Ratings Child History Photo LLaTAGCN 1.593 1.974 11.872 7.081 6.850 20.136 14.583 14.016 25.394 29.373 LLaTAGAT 2.592 2.974 12.871 8.080 7.849 21.138 15.583 15.014 26.393 30.369 LLaTASAGE 3.594 3.975 13.875 9.085 8.852 22.139 16.588 16.016 27.398 31.341
In this section, we present the complete runtime performance of LLaTA across 10 tagged datasets in Table 8, along with a comprehensive analysis of our method’s computational efficiency advantages.
Current graph structure learning (GSL) methods face significant efficiency bottlenecks, particularly in large-scale or noisy graphs. Traditional approaches—including metric learning, probabilistic modeling, and direct optimization—often require computationally expensive pairwise similarity calculations (e.g., kernel functions or attention mechanisms) or iterative optimization of adjacency matrices[36]. These operations exhibit quadratic or higher complexity relative to node counts, limiting scalability. For LLM-based GSL Methods (e.g., GraphEdit [7], LLM4RGNN [31]), efficiency issues are further exacerbated by the inherent overhead of large-model inference and fine-tuning. Empirical results reveal that these methods exceed 72 hours on larger datasets among the 10 benchmarks (As shown in Table 1), primarily due to their reliance on full-parameter tuning and dense graph operations.
Compared to existing paradigms, our framework achieves superior efficiency by: (a) Topology-Aware Context Enhancement: The Community-of-Thought mechanism provides high-quality topological context to guide LLM reasoning, eliminating the need for fine-tuning. By leveraging the inherent structure-awareness of communities, we reduce the dependency on costly end-to-end training while improving inference accuracy. (b) Parameter-Free Structure Reconstruction: A tree-based sampler enables lightweight graph restructuring without trainable parameters. This avoids the computational overhead of gradient-based adjacency matrix optimization (e.g., nuclear norm regularization in ProGNN [11]) or probabilistic sampling (e.g., Gumbel-Softmax in NeuralSparse [33]).
Enabled by these two key innovations, our framework achieves superior efficiency by: (1) Eliminating Fine-Tuning: Bypassing LLM parameter updates reduces GPU memory and time costs by orders of magnitude. (2) Sublinear Complexity: The tree sampler operates locally within communities, avoiding global pairwise calculations. (3) Parallelizability: Community-level operations enable distributed processing, scaling to graphs with $10^{5+}$ nodes where baseline LLM-GSL methods fail.